A flexible tool to experiment with distributed symbolic computation

Stéphane Dalmas and Marc Gaëtano ^{∗}

INRIA Sophia Antipolis, projet SAFIR 2004 route des lucioles BP 93 06902 Sophia-Antipolis CEDEX France

E-mail: stephane.dalmas@sophia.inria.fr

marc.gaetano@sophia.inria.fr

April 26, 1995

Abstract

In this paper we describe the Central Control and how it can be applied to easily make various experiments in distributed computation. The Central Control is a software component that has been designed to be the kernel of environments for scientific computations. Such an environment can offer a common and concurrent access to many tools needed by the scientist and the engineer: general purpose and specialized computer algebra systems, visualization tools, links with numerical libraries and tools to manipulate numerical programs etc.

The current version of the Central Control is a Scheme interpreter extended with new primitive types and procedures. To insure flexibility and to ease experimentations, there are only a few new essential procedures to provide the basic facilities (starting or connecting to servers, sending and receiving computations) together with a Scheme library that implements higher-level operations.

We will first describe the main features of the Central Control. We will then give examples of how it can be applied to problems that need or can benefit from distribution by showing examples of simple Scheme functions.

∗also I3S, CNRS URA 1376,Université de Nice Sophia–Antipolis